equation of locus of the point which divide double ordinates ordinates a of the ellipse x^2/a^2+y^2/b^2=1 inthe ratio 1:2 internally is

Grade:12

1 Answers

Shashank Joshi
7 years ago
Locus will be x^2/a^2+ 9y^2/b^2=1  ,   x coordinate of any point lying in that  double ordinate will be a\cos \theta
and y cordinate will  be get from section formula . which y= (b\sin \theta)/3 . noW JUST SQUARE AND ADD